OpenStudy (anonymous):

A car makes a trip due north for three-fourths of the time and due south one-fourth of the time. The average northward velocity has a magnitude of 42 m/s, and the average southward velocity has a magnitude of 46 m/s. Taking northward to be the positive direction, what is the average velocity for the trip?

OpenStudy (anonymous):

I'm not sure if this is correct, so someone correct me if I am wrong. .75(42m/s) - .25(46m/s) = 31.5 - 11.5 = 20 m/s

OpenStudy (anonymous):

it was correct!